# On estimation and order selection for multivariate extremes via clustering

### v2

@@ 3.2.1 Some deterministic estimates
Proposition 3.7.
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$ form an $m$-clustering of $W_{n}$ as defined in Definition 2.3 with respect to a dissimilarity measure $D$ defined in Definition 2.1. If $m<k$, then almost surely,
@@
By Proposition 2.5 with $S=B_{D}(\mathbf{a}_{i},\epsilon)$ and the choice $\epsilon<r_{A}$, we have each $H_{n}(S)=|W_{n}\cap B_{D}(\mathbf{a}_{i},\epsilon)|/|W_{n}|$ converges almost surely to $H(S)=p_{i}$, $i\in\{1,\ldots,k\}$. Hence, with probability $1$, the event $E_{n}(\epsilon,\delta)$ happens eventually as $n\rightarrow\infty$, namely, $\Pr\left(\liminf_{n}\mathbf{1}\left\{E_{n}(\epsilon,\delta)\right\}=1\right)=1$. Since $W_{n}$ satisfies the condition $\mathcal{A}(\epsilon,\delta)$ on $E_{n}(\epsilon,\delta)$, by Lemmas 3.3 and 3.6, for almost every outcome $\omega$ in the sample space $\Omega$, when $n$ is sufficiently large, we have when $m<k$ that
@@
The desired results follow if one takes $\limsup_{n}$ and $\liminf_{n}$ respectively in the two inequalities above, and then lets $\delta,\epsilon\rightarrow 0$ (see also Remark 3.2). ∎
@@
Proposition 3.8.
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$ form an $m$-clustering of $W_{n}$ as defined in Definition 2.3 with respect to a dissimilarity measure $D$ defined in Definition 2.1. Suppose $t>0$. If $m>k$, we have almost surely
@@
Proof.
The argument is similar to that of Proposition 3.7. In particular, under the restriction to the event $E_{n}(\epsilon,\delta)$ in (34), we have by Lemma 3.5 that for $m>k$
@@
Now we are ready to prove Theorem 3.1.
Proof of Theorem 3.1.
Putting together Propositions 3.7 and 3.8, and using the facts that $\bar{S}\in[0,1]$ and $P_{t}\in[0,1]$, we have almost surely that
@@ 4 Large deviation analysis of clustering-based spectral estimation
In this section, we provide a quantitative assessment of the consistency result in Corollary 2.6 through large-deviation-type bounds. This analysis is made possible through certain estimates used in the proof of Theorem 3.1 (see Section 3.2).
@@
Let $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$ be as defined in (18). Let $\left({A}_{k,n}=\left({\mathbf{a}}_{1,n}^{k},\ldots,{\mathbf{a}}_{k,n}^{k}\right),\mathfrak{C}_{k,n}=\left\{{C}_{1,n}^{k},\ldots,{C}_{k,n}^{k}\right\}\right)$ form a $k$-clustering of the extremal subsample $W_{n}$ as in (14). By Corollary 2.6, there exists permutation $\pi$, such that $\mathbf{a}_{\pi_{n}(i),n}^{k}$ and $p_{\pi_{n}(i),n}^{k}$ in (19) are consistent estimators for $\mathbf{a}_{i}$ and $p_{i}$, respectively. Note that an accurate estimation can be interpreted as that for small $x,y>0$, $D(\mathbf{a}_{\pi(i),n}^{k},\mathbf{a}_{i})<x$ and $|p_{\pi(i),n}^{k}-p_{i}|<y$ for all $i\in\{1,\ldots,k\}$. Now consider the complement “large deviation” event
@@
Proposition 4.3.
Suppose $\mathbf{X}$ satisfying (1) and (2) has a spectral measure of the following form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left({A}_{k,n}=\left({\mathbf{a}}_{1,n}^{k},\ldots,{\mathbf{a}}_{k,n}^{k}\right),\mathfrak{C}_{k,n}=\left\{{C}_{1,n}^{k},\ldots,{C}_{k,n}^{k}\right\}\right)$ form an $k$-clustering of $W_{n}$ as defined in Definition 2.3 with respect to a dissimilarity measure $D$ defined in Definition 2.1. Let $E(x,y)$ be the event defined in 37. Then for any $x,y>0$,
@@
Proof.
If $H_{n}\left(B_{D}(\mathbf{a}_{i},\epsilon)\right)=|W_{n}\cap B_{D}(\mathbf{a}_{i},\epsilon)|/|W_{n}|\geq p_{i}-\delta$ for all $i\in\{1,\ldots,k\}$, by Lemmas 3.4 and 3.6, as long as (31) holds, there exists a permutation $\pi:\{1,\ldots,k\}\mapsto\{1,\ldots,k\}$, such that $D(\mathbf{a}_{\pi(i),n}^{k},\mathbf{a}_{i})<\epsilon^{\prime}$ and $|p_{\pi(i),n}^{k}-p_{i}|\leq c_{k}\delta$ for all $i\in\{1,\ldots,k\}$. Hence under (31), whenever $\epsilon^{\prime}\leq x$ or $c_{k}\delta\leq y$,
@@
where $N$ and $B_{j}$’s are as in Lemma 4.1 with respective parameters $q_{1}$ and $q_{2}$ given as follows:
@@
as $n\rightarrow\infty$. Now applying Lemma 4.1, we have
@@
Remark 4.4.
The large-deviation-type estimates in Proposition 4.3 say that the probability $\Pr\left(E(x,y)\right)$ decays exponentially in the expected extremal subsample size $c_{(r)}\ell_{n}$. It is worth observing that the expression of $\Delta(x,y)$ reflects the following: The difficulty of clustering-based estimation measured by the aforementioned large error probabilities depends negatively on $p_{min}$ and $r_{A}$ (note that $\epsilon_{0}$ is an increasing function of $r_{A}$) and positively on $k$. In other words, the estimation is more accurate when the true discrete spectral measure has fewer atoms, the dissimilarity between the atoms is large, or the probability mass concentrated on each atom is significant.
We also have the following result which states that in the context of Theorem 3.1, the probability of false order election tends to 00 exponentially fast.
Proposition 4.5.
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, $k\in\mathbb{Z}_{+}$ where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$, $m\in\mathbb{Z}_{+}$, form an $m$-clustering of $W_{n}$ as defined in Definition 2.3 with respect to a dissimilarity measure $D$ defined in Definition 2.1. Let $r_{A}$ be defined as in (20), $p_{\min}$ be defined as in (24) and $t_{0}$ be defined as in (25). Then fix $t\in(0,t_{0})$,
@@
where $E_{n}(\epsilon,\delta)$ is in (34). Combining the inequalities regarding $\bar{S}$ in the proof of Proposition 3.7, and the inequalities regarding $P_{t}$ in the proof of Proposition 3.8, the event in the first probability on the right-hand side above is empty as long as $\delta>0$ satisfies
@@ 5.1 The models
where $\odot$ is interpreted as the matrix product with the sum operation replaced by the maximum operation. Note that due to the exchangeability of $\left(Z_{1},\ldots,Z_{k}\right)$, either model is identifiable only up to a permutation of the vectors $\mathbf{b}_{j}$, $j\in\{1,\ldots,k\}$, i.e. the distribution of $\mathbf{X}$ is unchanged if $B$ is replaced by $B_{\pi}:=\left(\mathbf{b}_{\pi(1)},\ldots,\mathbf{b}_{\pi(k)}\right)$ for any permutation $\pi:\{1,\ldots,k\}\mapsto\{1,\ldots,k\}$. The models of types (41) and (42) have recently attracted considerable interest in connection with causal structural equations for extremes; see, e.g., [14, 15].
@@ 5.2 Order selection and coefficient estimation
Due to the discrete nature of the spectral measure, the likelihood functions of these models are inaccessible (see, e.g., [7, 34, 8]). Even without taking a perspective of extremes, the max-linear model (42) does not admit a smooth density. Therefore, the usual model selection techniques based on information criteria are not available. On the other hand, the spectral measure of these factor models, including (39) and (40), is of the form (18). Therefore, the penalized ASW method proposed in Section 3 could be used to select the order of factors $k$, whose consistency is supported by Theorem 3.1.
Suppose from now on the order $k$ is assumed to be known. Another noteworthy issue deserving discussion is whether we can translate the estimation of the spectral measure through a $k$-clustering algorithm (refer to Section 2.3) into an estimation of the coefficient matrix $B=\left(\mathbf{b}_{1},\ldots,\mathbf{b}_{k}\right)$ in (41) or (42). Note that the constraint (44) also needs to be taken into account. Combining (43) and (44), to solve the $kd$ coefficients in $B$ from $p_{j}$’s and $\mathbf{a}_{j}$’s, we have totally $kd+d-1$ free equations ($k-1$ from the equations for $p_{j}$’s, $(d-1)k$ from the equations for $\mathbf{a}_{j}$’s and $d$ from (44)). When $p_{j}$’s and $\mathbf{a}_{j}$’s are estimated via $k$-clustering, the over-determined system may not admit a solution, although this over-determined relation holds asymptotically in view of Corollary 2.6.
@@
Note that if $\|\cdot\|_{(s)}=\|\cdot\|_{\alpha}$ already, then $\|\mathbf{a}_{j}\|_{\alpha}=1$. So one can plug in estimated $\mathbf{a}_{j}$ and $p_{j}$ via $k$-clustering on the $\alpha$-norm sphere into (45), obtaining, say, $\widehat{\mathbf{b}}_{j}$, $j\in\{1,\ldots,k\}$. However, the condition (44) may not be satisfied. We propose the following simple correction: first, form the preliminary estimated coefficient matrix $\widehat{B}:=\left(\widehat{\mathbf{b}}_{1},\ldots,\widehat{\mathbf{b}}_{k}\right)=:\left({\mathbf{r}}_{1},\ldots,{\mathbf{r}}_{d}\right)^{\top}$, where $\mathbf{r}_{i}^{\top}$, $i\in\left\{1,\ldots,d\right\}$, are row vectors of $\widehat{B}$. Then we obtain the final estimate $\widetilde{B}=\left(\widetilde{\mathbf{b}}_{1},\ldots,\widetilde{\mathbf{b}}_{k}\right)$ of $B$ through replacing each row $\mathbf{r}_{i}$ by $\mathbf{r}_{i}/\|\mathbf{r}_{i}\|_{\alpha}$, which ensures (44). It follows from Corollary 2.6 and a continuous mapping argument that the thus obtained estimate of $B$ is consistent (up to a permutation of $\mathbf{b}_{i}$’s).

### v3

@@ 3.2.2 Proof of Theorem 1
Proposition 2.
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$ form an $m$-clustering of $W_{n}$ as defined in Definition 2 with respect to a dissimilarity measure $D$ defined in Definition 1. If $m<k$, then almost surely,
@@
By Proposition 1 with $S=B_{D}(\mathbf{a}_{i},\epsilon)$ and the choice $\epsilon<r_{A}$, we have each $H_{n}(S)=|W_{n}\cap B_{D}(\mathbf{a}_{i},\epsilon)|/|W_{n}|$ converges almost surely to $H(S)=p_{i}$, $i\in\{1,\ldots,k\}$. Hence, with probability $1$, the event $E_{n}(\epsilon,\delta)$ happens eventually as $n\rightarrow\infty$, namely, $\Pr\left(\liminf_{n}\mathbf{1}\left\{E_{n}(\epsilon,\delta)\right\}=1\right)=1$. Since $W_{n}$ satisfies the condition $\mathcal{A}(\epsilon,\delta)$ on $E_{n}(\epsilon,\delta)$, by Lemmas 1 and 4, for almost every outcome $\omega$ in the sample space $\Omega$, when $n$ is sufficiently large, we have when $m<k$ that
@@
The desired results follow if one takes $\limsup_{n}$ and $\liminf_{n}$ respectively in the two inequalities above, and then lets $\delta,\epsilon\rightarrow 0$ (see also Remark 4). ∎
@@
Proposition 3.
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$ form an $m$-clustering of $W_{n}$ as defined in Definition 2 with respect to a dissimilarity measure $D$ defined in Definition 1. Suppose $t>0$. If $m>k$, we have almost surely
@@
Proof:.
The argument is similar to that of Proposition 2. In particular, under the restriction to the event $E_{n}(\epsilon,\delta)$ in (34), we have by Lemma 3 that for $m>k$
@@
Now we are ready to prove Theorem 1.
Proof of Theorem 1:.
Putting together Propositions 2 and 3, and using the facts that $\bar{S}\in[0,1]$ and $P_{t}\in[0,1]$, we have almost surely that
@@ 4 Large deviation analysis of clustering-based spectral estimation
In this section, we provide a quantitative assessment of the consistency result in Corollary 1 through large-deviation-type bounds. This analysis is made possible through certain estimates used in the proof of Theorem 1 (see Section 3.2).
@@
Let $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$ be as defined in (18). Let $\left({A}_{k,n}=\left({\mathbf{a}}_{1,n}^{k},\ldots,{\mathbf{a}}_{k,n}^{k}\right),\mathfrak{C}_{k,n}=\left\{{C}_{1,n}^{k},\ldots,{C}_{k,n}^{k}\right\}\right)$ form a $k$-clustering of the extremal subsample $W_{n}$ as in (14). By Corollary 1, there exists permutation $\pi$, such that $\mathbf{a}_{\pi_{n}(i),n}^{k}$ and $p_{\pi_{n}(i),n}^{k}$ in (19) are consistent estimators for $\mathbf{a}_{i}$ and $p_{i}$, respectively. Note that an accurate estimation can be interpreted as that for small $x,y>0$, $D(\mathbf{a}_{\pi(i),n}^{k},\mathbf{a}_{i})<x$ and $|p_{\pi(i),n}^{k}-p_{i}|<y$ for all $i\in\{1,\ldots,k\}$. Now consider the complement “large deviation” event
@@
Proposition 4.
Suppose $\mathbf{X}$ satisfying (1) and (2) has a spectral measure of the following form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left({A}_{k,n}=\left({\mathbf{a}}_{1,n}^{k},\ldots,{\mathbf{a}}_{k,n}^{k}\right),\mathfrak{C}_{k,n}=\left\{{C}_{1,n}^{k},\ldots,{C}_{k,n}^{k}\right\}\right)$ form an $k$-clustering of $W_{n}$ as defined in Definition 2 with respect to a dissimilarity measure $D$ defined in Definition 1. Let $E(x,y)$ be the event defined in 37. Then for any $x,y>0$,
@@
Proof:.
If $H_{n}\left(B_{D}(\mathbf{a}_{i},\epsilon)\right)=|W_{n}\cap B_{D}(\mathbf{a}_{i},\epsilon)|/|W_{n}|\geq p_{i}-\delta$ for all $i\in\{1,\ldots,k\}$, by Lemmas 2 and 4, as long as (31) holds, there exists a permutation $\pi:\{1,\ldots,k\}\mapsto\{1,\ldots,k\}$, such that $D(\mathbf{a}_{\pi(i),n}^{k},\mathbf{a}_{i})<\epsilon^{\prime}$ and $|p_{\pi(i),n}^{k}-p_{i}|\leq c_{k}\delta$ for all $i\in\{1,\ldots,k\}$. Hence under (31), whenever $\epsilon^{\prime}\leq x$ or $c_{k}\delta\leq y$,
@@
where $N$ and $B_{j}$’s are as in Lemma 5 with respective parameters $q_{1}$ and $q_{2}$ given as follows:
@@
as $n\rightarrow\infty$. Now applying Lemma 5, we have
@@
Remark 6.
The large-deviation-type estimates in Proposition 4 say that the probability $\Pr\left(E(x,y)\right)$ decays exponentially in the expected extremal subsample size $c_{(r)}\ell_{n}$. It is worth observing that the expression of $\Delta(x,y)$ reflects the following: The difficulty of clustering-based estimation measured by the aforementioned large error probabilities depends negatively on $p_{min}$ and $r_{A}$ (note that $\epsilon_{0}$ is an increasing function of $r_{A}$) and positively on $k$. In other words, the estimation is more accurate when the true discrete spectral measure has fewer atoms, the dissimilarity between the atoms is large, or the probability mass concentrated on each atom is significant.
We also have the following result which states that in the context of Theorem 1, the probability of false order election tends to 00 exponentially fast.
Proposition 5.
Suppose $\mathbf{X}$ satisfying (1) and (2) has a discrete spectral measure of the form $H=\sum_{i=1}^{k}p_{i}\delta_{\mathbf{a}_{i}}$, $k\in\mathbb{Z}_{+}$ where $\mathbf{a}_{i}$’s are distinct points on $\mathbb{S}_{+}^{d-1}$, and $p_{i}>0$, $p_{1}+\ldots+p_{k}=1$. Let $W_{n}$ denote the extremal subsample as in (14), and $\left(A_{m,n},\mathfrak{C}_{m,n}\right)$, $m\in\mathbb{Z}_{+}$, form an $m$-clustering of $W_{n}$ as defined in Definition 2 with respect to a dissimilarity measure $D$ defined in Definition 1. Let $r_{A}$ be defined as in (20), $p_{\min}$ be defined as in (24) and $t_{0}$ be defined as in (25). Then fix $t\in(0,t_{0})$,
@@
where $E_{n}(\epsilon,\delta)$ is in (34). Combining the inequalities regarding $\bar{S}$ in the proof of Proposition 2, and the inequalities regarding $P_{t}$ in the proof of Proposition 3, the event in the first probability on the right-hand side above is empty as long as $\delta>0$ satisfies
@@ 5.1 The models
where $\odot$ is interpreted as the matrix product with the sum operation replaced by the maximum operation. Note that due to the exchangeability of $\left(Z_{1},\ldots,Z_{k}\right)$, either model is identifiable only up to a permutation of the vectors $\mathbf{b}_{j}$, $j\in\{1,\ldots,k\}$, i.e., the distribution of $\mathbf{X}$ is unchanged if $B$ is replaced by $B_{\pi}:=\left(\mathbf{b}_{\pi(1)},\ldots,\mathbf{b}_{\pi(k)}\right)$ for any permutation $\pi:\{1,\ldots,k\}\mapsto\{1,\ldots,k\}$. The models of types (41) and (42) have recently attracted considerable interest in connection with causal structural equations for extremes; see, e.g., [14, 15].
@@ 5.2 Order selection and coefficient estimation
Due to the discrete nature of the spectral measure, the likelihood functions of these models are inaccessible (see, e.g., [7, 34, 8]). Even without taking a perspective of extremes, the max-linear model (42) does not admit a smooth density. Therefore, the usual model selection techniques based on information criteria are not available. On the other hand, the spectral measure of these factor models, including (39) and (40), is of the form (18). Therefore, the penalized ASW method proposed in Section 3 could be used to select the order of factors $k$, whose consistency is supported by Theorem 1.
Suppose from now on the order $k$ is assumed to be known. Another noteworthy issue deserving discussion is whether we can translate the estimation of the spectral measure through a $k$-clustering algorithm (refer to Section 2.3) into an estimation of the coefficient matrix $B=\left(\mathbf{b}_{1},\ldots,\mathbf{b}_{k}\right)$ in (41) or (42). Note that the constraint (44) also needs to be taken into account. Combining (43) and (44), to solve the $kd$ coefficients in $B$ from $p_{j}$’s and $\mathbf{a}_{j}$’s, we have totally $kd+d-1$ free equations ($k-1$ from the equations for $p_{j}$’s, $(d-1)k$ from the equations for $\mathbf{a}_{j}$’s and $d$ from (44)). When $p_{j}$’s and $\mathbf{a}_{j}$’s are estimated via $k$-clustering, the over-determined system may not admit a solution, although this over-determined relation holds asymptotically in view of Corollary 1.
@@
Note that if $\|\cdot\|_{(s)}=\|\cdot\|_{\alpha}$ already, then $\|\mathbf{a}_{j}\|_{\alpha}=1$. So one can plug in estimated $\mathbf{a}_{j}$ and $p_{j}$ via $k$-clustering on the $\alpha$-norm sphere into (45), obtaining, say, $\widehat{\mathbf{b}}_{j}$, $j\in\{1,\ldots,k\}$. However, the condition (44) may not be satisfied. We propose the following simple correction: first, form the preliminary estimated coefficient matrix $\widehat{B}:=\left(\widehat{\mathbf{b}}_{1},\ldots,\widehat{\mathbf{b}}_{k}\right)=:\left({\mathbf{r}}_{1},\ldots,{\mathbf{r}}_{d}\right)^{\top}$, where $\mathbf{r}_{i}^{\top}$, $i\in\left\{1,\ldots,d\right\}$, are row vectors of $\widehat{B}$. Then we obtain the final estimate $\widetilde{B}=\left(\widetilde{\mathbf{b}}_{1},\ldots,\widetilde{\mathbf{b}}_{k}\right)$ of $B$ through replacing each row $\mathbf{r}_{i}$ by $\mathbf{r}_{i}/\|\mathbf{r}_{i}\|_{\alpha}$, which ensures (44). It follows from Corollary 1 and a continuous mapping argument that the thus obtained estimate of $B$ is consistent (up to a permutation of $\mathbf{b}_{i}$’s).